# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v1

@@ Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect their reasoning capabilities remains poorly understood. We investigate whether multi-agent systems function as strategic reasoners capable of payoff optimization or as identity-driven actors that prioritize role alignment over explicit incentives. We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities. We do so through systematic experiments with four LLM architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games. We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e. when personas are active, all models achieve near-zero Nash equilibrium rates (0–6.7%) in scenarios where payoff-optimal equilibria exist, despite having complete payoff information. Removing personas and providing explicit payoffs enables Qwen models to achieve high Nash equilibrium rates (65–90%) in these scenarios, demonstrating that both conditions are necessary for strategic reasoning capabilities. Personas systematically bias equilibrium selection toward socially preferred outcomes, with 100% of Nash equilibria being Green Transition (socially preferred) when personas are present, while models completely fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, revealing that design choices about personas and payoff visibility interact in ways that fundamentally determine system behavior. We observe striking model-dependent patterns where Qwen architectures show high sensitivity to both personas and payoff visibility, while Llama and Mistral maintain rigid reasoning patterns across conditions. These findings establish that representational choices are substantive governance decisions that determine whether multi-agent systems function as strategic reasoners or identity-driven actors, with important implications for deployment in strategic settings.
@@ 1. Introduction
This question determines whether multi-agent LLM systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025). If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024). This creates a fundamental tension between the representational goals of persona-based systems and their ability to reason strategically when payoff information is available (Bender et al., 2021; Duan et al., 2024b). Understanding how personas interact with payoff information in multi-agent settings is essential to make deployment decisions (Weidinger et al., 2022b; Chan et al., 2024; Casper et al., 2024; Rivera et al., 2024; Chan et al., 2024). However, this gap remains largely unexplored. Existing work has examined personas, payoff visibility, and model differences separately but has not systematically studied how these factors together determine system behavior in strategic settings (Neumann et al., 2025; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024).
@@
Understanding how role identity and payoff optimization interact is critical in multi-agent strategic games. We study environmental policy scenarios with four agents (Industrialist, Government, Environmental Activist, and Citizen) coordinating to reach Nash equilibria (Nash, 1951), which can yield either Green Transition (socially preferred and environmentally sustainable) or Tragedy of Commons (payoff-optimal but environmentally damaging) outcomes (Hardin, 1968). Through a $\times2$ factorial experiment manipulating persona presence and payoff visibility across four model architectures, we examine whether agents prioritize role-aligned actions or payoff-optimal strategies. Our findings show that personas fundamentally alter strategic reasoning, preventing payoff optimization: when personas are active, all models achieve near-zero Nash equilibrium rates in economic scenarios where Tragedy of Commons is payoff-optimal, despite complete payoff information. Removing personas and providing explicit payoffs enables certain models to achieve high Nash equilibrium in these scenarios, demonstrating both conditions are necessary for strategic reasoning. Personas systematically bias equilibrium selection toward socially preferred outcomes, with Nash equilibria being exclusively Green Transition when personas are present, while models fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, and we observe model-dependent patterns where some architectures show high sensitivity to both personas and payoff visibility, while others maintain rigid reasoning patterns across conditions.
This work provides key contributions to understanding LLM behavior in multi-agent strategic settings. We demonstrate that role identity bias completely overrides explicit payoff optimization, preventing Nash equilibrium. Additionally, personas and payoff visibility interact in ways that fundamentally determine system behavior, meaning design choices about representation formats are substantive governance decisions. We also reveal model-dependent persona effects, suggesting different architectures implement strategic reasoning through different mechanisms. These findings have practical implications for deployment, as personas alter system behavior in ways that may not be immediately apparent to users, showing that deploying personas requires careful consideration of how identity interacts with payoff optimization in strategic decision making.
@@ 2.1. LLMs in Multi-Agent Systems
Large language models (LLMs) have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996). Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024). Following studies explored deliberative and collaborative agent frameworks, emphasizing consensus formation and group-level reasoning (Abdelnabi et al., 2024; Bianchi et al., 2024b; Agashe et al., 2025; Bo et al., 2024).
@@
However, most of this literature focuses on task-oriented collaboration or collective sensemaking, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
@@ 3.1.2. Scenarios
We tested 53 scenarios across three scenario types, all designed with payoff-coupled narratives where the scenario description aligns with the underlying payoff structure. All scenario narratives were generated using GPT-5-mini [^1] Azure OpenAI GPT-5-mini (model version 2025-08-07) and subsequently reviewed and filtered to ensure consistency with the intended payoff structure and equilibrium class [^2] Model-generated scenarios were used to reduce author-specific framing bias and to enable systematic variation across narrative surface forms while holding the underlying game structure fixed. Scenarios were retained only if the narrative description unambiguously aligned with the corresponding payoff incentives, and no model was evaluated on scenarios it had previously generated.. Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided. Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix D.:
@@
Table 2. Environmental Scenario Nash Equilibrium Rates: With Personas vs Without Personas. This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
@@
3.2. Chain-of-Thought Reasoning Analysis
We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes. For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix F for keyword categories). We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style. From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making). This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies. While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here. A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented in the Appendix (Figures 5-8).
3.3. Models and Experimental Conditions
@@
4. Results
We now present quantitative results organized by the research questions that we aim to answer. We report results by answering the following research questions.
Answer: Yes. When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning.
Fig. 2 shows our findings. When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information. Models select Green Transition actions (20-59% [^5] Fig. 4 in Appendix shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
@@ 4. Results
Figure 2. Economic Scenarios Nash Equilibrium Rates by Model and Condition. This heatmap visualizes Nash equilibrium achievement rates in economic scenarios (where Tragedy of Commons is the payoff-optimal equilibrium) across all four models and four experimental conditions. The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
@@
Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions. When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models. This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices. CoT analysis confirms this claim. With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions. This pattern reverses when personas are removed. Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference). CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization. This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed. However, this capability is model-dependent. Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
@@
Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help
@@
Table 6(a) quantifies the magnitude of persona effects on overall Nash equilibrium rates. All models show negative persona effects in the visible payoffs variant (-5 to -31 pp), but the mechanisms differ. For Qwen models, removing personas enables payoff-optimal Tragedy equilibrium in economic scenarios but reduces Green Transition equilibrium in environmental scenarios when visible payoffs are present (96.1% $\to$ 16.6% Nash for Qwen-32B, Table 5), resulting in net negative overall effects (-20-25 pp, Table 6(a)). For Llama/Mistral, removing personas only reduces environmental Nash equilibrium without enabling economic Nash (0% remains), reflecting pure Nash rate loss. Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
Analysing the Chain-Of-Thought of the models shows how differently they reason. Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence. Mistral shows payoff optimization keywords but cannot identify Tragedy equilibrium regardless of condition. Qwen models show dramatic shifts with and without personas. With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning. 99.7% of the rationals [^7] See Section H for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning. Additionally, model size does not predict persona dependence. Both Qwen-7B and Qwen-32B show similar reasoning pattern shifts, but Qwen-32B achieves higher Nash rates due to stronger payoff optimization.
@@
Answer: Yes. When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased.
Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants. Fig. 3 reveals model-dependent patterns when personas are removed. For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition. For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 2 means most economic Nash is Tragedy. For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
Table 6. Persona Effects on Nash Equilibrium Rates: Overall (left) and Economic Scenarios (right). Left table shows overall persona effects across all scenarios (economic and environmental combined). Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal). Negative values indicate personas help and positive values indicate personas hurt.
(a) Overall Nash Equilibrium Rates

### v2

@@ Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect their reasoning capabilities remains poorly understood. We investigate whether multi-agent systems function as strategic reasoners capable of payoff optimization or as identity-driven actors that prioritize role alignment over explicit incentives. We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities. We do so through systematic experiments with four Large language model (LLM) architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games. We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e. when personas are active, all models achieve near-zero Nash equilibrium rates (0–6.7%) in scenarios where payoff-optimal equilibria exist, despite having complete payoff information. Removing personas and providing explicit payoffs enables Qwen models to achieve high Nash equilibrium rates (65–90%) in these scenarios, demonstrating that both conditions are necessary for strategic reasoning capabilities. Personas systematically bias equilibrium selection toward socially preferred outcomes, with 100% of Nash equilibria being Green Transition (socially preferred) when personas are present, while models completely fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, revealing that design choices about personas and payoff visibility interact in ways that fundamentally determine system behavior. We observe striking model-dependent patterns where Qwen architectures show high sensitivity to both personas and payoff visibility, while Llama and Mistral maintain rigid reasoning patterns across conditions. These findings establish that representational choices are substantive governance decisions that determine whether multi-agent systems function as strategic reasoners or identity-driven actors, with important implications for deployment in strategic settings.
@@ 1. Introduction
This question determines whether multi-agent large language model (LLM) systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025). If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024). This creates a fundamental tension between the representational goals of persona-based systems and their ability to reason strategically when payoff information is available (Bender et al., 2021; Duan et al., 2024b). Understanding how personas interact with payoff information in multi-agent settings is essential to make deployment decisions (Weidinger et al., 2022b; Chan et al., 2024; Casper et al., 2024; Rivera et al., 2024; Chan et al., 2024). However, this gap remains largely unexplored. Existing work has examined personas, payoff visibility, and model differences separately but has not systematically studied how these factors together determine system behavior in strategic settings (Neumann et al., 2025; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024).
@@
Understanding how role identity and payoff optimization interact is critical in multi-agent strategic games. We study environmental policy scenarios with four agents (Industrialist, Government, Environmental Activist, and Citizen) coordinating to reach Nash equilibria (Nash, 1951), which can yield either Green Transition (socially preferred and environmentally sustainable) or Tragedy of Commons (payoff-optimal but environmentally damaging) outcomes (Hardin, 1968). Through a $\times2$ factorial experiment manipulating persona presence and payoff visibility across four model architectures, we examine whether agents prioritize role-aligned actions or payoff-optimal strategies. Our findings show that personas fundamentally alter strategic reasoning, preventing payoff optimization: when personas are active, all models achieve near-zero Nash equilibrium rates in economic scenarios where Tragedy of Commons is payoff-optimal, despite complete payoff information. Removing personas and providing explicit payoffs enables certain models to achieve high Nash equilibrium in these scenarios, demonstrating both conditions are necessary for strategic reasoning. Personas systematically bias equilibrium selection toward socially preferred outcomes, with Nash equilibria being exclusively Green Transition when personas are present, while models fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, and we observe model-dependent patterns where some architectures show high sensitivity to both personas and payoff visibility, while others maintain rigid reasoning patterns across conditions. This work provides key contributions to understanding LLM behavior in multi-agent strategic settings. We demonstrate that role identity bias completely overrides explicit payoff optimization, preventing Nash equilibrium. Additionally, personas and payoff visibility interact in ways that fundamentally determine system behavior, meaning design choices about representation formats are substantive governance decisions. We also reveal model-dependent persona effects, suggesting different architectures implement strategic reasoning through different mechanisms. These findings have practical implications for deployment, as personas alter system behavior in ways that may not be immediately apparent to users, showing that deploying personas requires careful consideration of how identity interacts with payoff optimization in strategic decision making.
@@ 2.1. LLMs in Multi-Agent Systems
Large language models have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996). Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024). Following studies explored deliberative and collaborative agent frameworks, emphasizing consensus formation and group-level reasoning (Abdelnabi et al., 2024; Bianchi et al., 2024b; Agashe et al., 2025; Bo et al., 2024).
@@
However, most of this literature focuses on task-oriented collaboration or collective sense-making, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
@@ 3.1.2. Scenarios
We tested 53 scenarios across three scenario types, all designed with payoff-coupled narratives where the scenario description aligns with the underlying payoff structure. All scenario narratives were generated using GPT-5-mini [^1] Azure OpenAI GPT-5-mini (model version 2025-08-07) and subsequently reviewed and filtered to ensure consistency with the intended payoff structure and equilibrium class [^2] Model-generated scenarios were used to reduce author-specific framing bias and to enable systematic variation across narrative surface forms while holding the underlying game structure fixed. Scenarios were retained only if the narrative description unambiguously aligned with the corresponding payoff incentives, and no model was evaluated on scenarios it had previously generated.. Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided. Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix A.7.:
@@
3.2. Chain-of-Thought Reasoning Analysis
We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes. For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix C for keyword categories). We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style. From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making). This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies. While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here. A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented in the Appendix (Figures 5-8).
3.3. Models and Experimental Conditions
4. Results
We now present quantitative results organized by the research questions that we aim to answer. We report results by answering the five research questions ,across the settings mentioned above, based on our experiments backed by tables and figures.
Answer: Yes. When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning. Fig. 3 shows our findings. When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information. Models select Green Transition actions (20-65%, Fig. 2 shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
Table 2. Environmental Scenario Nash Equilibrium Rates: With Personas vs. Without Personas. This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
@@ 4. Results
Figure 2. Green Transition Action Selection Rates in Economic Scenarios. This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where the Tragedy of Commons is the payoff-optimal equilibrium.
Figure 3. Economic Scenarios Nash Equilibrium Rates by Model and Condition. This heatmap visualizes Nash equilibrium achievement rates in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium across all four models and four experimental conditions. The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
@@
Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions. When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models. This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices. CoT analysis confirms this claim. With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions.
This pattern reverses when personas are removed. Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference). CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization. This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed. However, this capability is model-dependent. Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
@@
Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs. Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help.
@@
Table 6(a) quantifies the magnitude of persona effects on overall Nash equilibrium rates. All models show negative persona effects in the visible payoffs variant (-5 to -31 pp), but the mechanisms differ.
For Qwen models, removing personas enables payoff-optimal Tragedy equilibrium in economic scenarios but reduces Green Transition equilibrium in environmental scenarios when visible payoffs are present (96.1% $\to$ 16.6% Nash for Qwen-32B, Table 5), resulting in net negative overall effects (-20-25 pp, Table 6(a)). For Llama/Mistral, removing personas only reduces environmental Nash equilibrium without enabling economic Nash (0% remains), reflecting pure Nash rate loss. Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
Analyzing the Chain-Of-Thought of the models shows how differently they reason. Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence. Mistral shows payoff optimization keywords but cannot identify Tragedy equilibrium regardless of condition. Qwen models show dramatic shifts with and without personas. With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning. 99.7% of the rationales [^6] See Section E for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning.
Additionally, model size does not predict persona dependence. Both Qwen-7B and Qwen-32B show similar reasoning pattern shifts, but Qwen-32B achieves higher Nash rates due to stronger payoff optimization.
@@
Answer: Yes. When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased. Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants. Fig. 4 reveals model-dependent patterns when personas are removed. For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition. For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 3 means most economic Nash is Tragedy. For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
Table 6. Persona Effects on Nash Equilibrium Rates: Overall (left) and Economic Scenarios (right). Left table shows overall persona effects across all scenarios (economic and environmental combined). Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal). Negative values (percentage points) indicate personas help and positive values (percentage points) indicate personas hurt.
(a) Overall Nash Equilibrium Rates and their dependence on personas for various LLM architectures.